OpenStudy (anonymous):

You toss a coin and roll a number cube. Find P(heads and an even number). how do i do this???

OpenStudy (anonymous):

you mean p is the probability?

OpenStudy (anonymous):

yes

OpenStudy (across):

\(P_{\text{coin}}(\text{heads})=0.5\) \(P_{\text{dice}}(\text{even})=0.5\)

OpenStudy (callisto):

sorry you need to multiply the 2 numbers as it's an 'and' case, P(head and even no) = 0.5x0.5 = 0.25

OpenStudy (apoorvk):

sample space={1T 2T 3T 4T 5T 6T 1H 2H 3H 4H 5H 6H } TOTAL NO. of outcomes=12 no. of favourable outcomes = 3 ... {2H 4H 6H} SO, p= 3/12 = 1/4

OpenStudy (anonymous):